The perimeter of a rectangular carpet is 70 feet. The width is three-fourth the length. Find the width IM KIND OF LOST!
Answer by mananth(16949)

The perimeter of a rectangular carpet is 70 feet. The width is three-fourth the length. Find the width IM KIND OF LOST!
let the length be x
the width is 3x/4
Perimeter of rectangle = 2L+2W
2x+ 2 * 3x/4 = 70
2x+3x/2 = 70
4x+3x / 2= 70
7x= 140
x= 20 feet which is the length
3/4 *20 = 15 feet is the width
Now dont get lost. cheers!